OpenStudy (anonymous):

all terms in a geometric sequence are positive integers, and the first 3 terms are n, n+3 and 2n+6. what is the 4th term of this sequence?

OpenStudy (anonymous):

in a geometric sequence the common ratio = 2nd/term / first ( ar / a) so here: common ratio = n+3 / n so 4th term = (2n + 6) * ( n+3) / n = (2n^2 + 12n + 18) / n = 2n + 12 + 18/n

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw